Perform the requested operation or operations. f(x) = 4x + 6, g(x) = 5x2
Find (f + g)(x).

Answer by jim_thompson5910(35256)

(f + g)(x) = f(x) + g(x)

(f + g)(x) = 4x + 6 + 5x^2

(f + g)(x) = 5x^2 + 4x + 6